# {centering} Standard Model anomalies and vacuum stability for lepton portals with extra $U(1)$ symmetry

### v1

@@ 1 Introduction
Recently, the measured values of the muon $(g-2)_{\mu}$ [4, 5] and the $W$ boson mass $m_{{}_{W}}$ [6] have both indicated significant deviations from the SM predictions, that we will refer to as the SM anomalies in the following, and have motivated the search for an explanation within extensions of the SM with extra particles and symmetries. In this context, it becomes of interest to investigate what the impact of new physics on the stability issue is in these phenomenologically motivated models, as discussed above.
In this article, we revisit a lepton portal model with an extra $U(1)^{\prime}$ gauge symmetry from this perspective, and perform an RG analysis to investigate on its stability. In this model, suggested by some of the authors [7], an $SU(2)_{L}$ singlet vector-like lepton charged under $U(1)^{\prime}$ is introduced. The latter mixes with the muon through the vacuum expectation values (VEVs) of an extra Higgs doublet and a dark Higgs field, leading to a small seesaw mass for the muon. The muon $g-2$ can be explained thanks to extra one-loop corrections coming from the vector-lepton and the $Z^{\prime}$ boson, while the $W$ mass $m_{{}_{W}}$ can be increased, at the same time, through a tree-level mixing between the $Z$ boson and the $Z^{\prime}$ boson when the $Z^{\prime}$ is heavier than $Z$. As the muon $g-2$ anomaly favors the $U(1)^{\prime}$ breaking scale to be below about $200\,{\rm GeV}$, the extra gauge coupling required to increase $m_{{}_{W}}$ tends to be large [7], so that it can have a strong impact on the RG flow of the theory.
@@ 3.1 SM anomalies: muon $g-2$ and $W$ boson mass
We comment on how flexible the parameter space is to explain the muon $g-2$ and $W$ boson mass anomalies. In the left plot of Fig. 1, we depict the one-loop correction to the muon $g-2$ as a function of $m_{{}_{Z^{\prime}}}$ in our model, with $v_{s}=v_{\phi}/\sqrt{2}=200\,{\rm GeV}$ and $150\,{\rm GeV}$. We set the parameters of the vector-like lepton sector to $M_{E}=1\,{\rm TeV}$, $\sin\theta_{L}=\sin\theta_{R}=0.01$, and choose $\sin\beta=0.18$ and $\lambda_{\phi}=0.1$ [7]. In the right plot of Fig. 1, we also show the tree-level correction to the $W$ boson mass due to the $Z$-$Z^{\prime}$ mixing as a function of $m_{{}_{Z^{\prime}}}$, with $v_{s}=v_{\phi}/\sqrt{2}=200\,{\rm GeV}$ and $150\,{\rm GeV}$, making the same choice for the other parameters as in the previous case.
@@ 3.2 $Z^{\prime}$ signatures at the LHC
Let us now discuss the $Z^{\prime}$ production and its decay to the dilepton state. The $Z^{\prime}$ production cross sections (left) and $Z^{\prime}$ production cross sections multiplied by the $ee$ and $\mu\mu$ dilepton branching ratio (right) for $pp$ collision at $13$ TeV are depicted in Fig. 3 as a function of $m_{{}_{Z}^{\prime}}$. The opening of the $WW$ and $hZ$ channels is such that the dilepton production is suppressed strongly for large enough values of $m_{{}_{Z^{\prime}}}$, for e.g. it is less than $10^{-1}$ for $m_{{}_{Z^{\prime}}}\gtrsim 300~{}$GeV.
With the results of ATLAS for dilepton final states at the $13$ TeV LHC run with integrated luminosity $L_{\mathrm{int}}=139$/fb [15], parametric regions with $m_{{}_{Z^{\prime}}}\gtrsim 300~{}$ GeV and $\sin\beta=0.18$ are allowed in our scenario. As we have seen, within this parametric region the $WW$ and $hZ$ final states are dominant, and these resonant boson productions are safe with respect to the $L_{\mathrm{int}}=139$/fb, $13$ TeV LHC results  [18, 16, 17, 19]. As the LHC constraints on dileptons are quite severe, dilepton final state produced from $Z^{\prime}$ boson resonance around $m_{{}_{Z^{\prime}}}\sim 300~{}$GeV should be targeted by the next LHC results. The same considerations hold for $WW$ final states for $m_{{}_{Z^{\prime}}}\gtrsim 300~{}$ GeV, whose dominant mode can be targeted with our scenario.
@@ 4.5 $Z^{\prime}$ mass dependence of the Landau pole
and calculating its Landau pole $\mu_{{}_{L}}=\mu_{{}_{\rm IR}}e^{{6\pi^{2}}/{7g_{{}_{Z^{\prime}}}(\mu_{{}_{\rm IR}})}}$ for the different IR values of $g_{{}_{Z^{\prime}}}$ in the two sets of the parameters in Table 4. The Landau pole of the full model is compared to $\mu_{{}_{L}}$ in Fig. 5, where it can be seen that the higher the value of $m_{{}_{Z^{\prime}}}$ ($g_{{}_{Z^{\prime}}}$), the closer the two are. It is easy to conclude from this plot that, for “high enough” (IR) values of $g_{{}_{Z^{\prime}}}$, the Landau pole is almost fully determined by the gauge coupling, that, through its contribution in their beta functions, causes the other couplings to diverge at slightly lower values. Lowering the (IR) value of $g_{{}_{Z^{\prime}}}$, the Landau pole of (45) becomes considerably higher than that of the full model. In this region of parameter space, the appearance of the Landau pole is a more “collective” phenomenon. All the couplings conspire together to generate it, and the determination of the pole scale from the RG equation of $g_{{}_{Z^{\prime}}}$ taken in isolation does not provide a good approximation to the pole scale of the full model any more.

### v3

@@ 1 Introduction
Recently, the measured values of the muon $(g-2)_{\mu}$ [4, 5] and the $W$ boson mass $m_{{}_{W}}$ [6] have both indicated significant deviations from the SM predictions [7], that we will refer to as the SM anomalies in the following, and have motivated the search for an explanation within extensions of the SM with extra particles and symmetries. However, it has been recently shown that the SM prediction for the muon $g-2$ based on the lattice results is consistent with its experimental value [8, 9]. Furthermore, the recent CMD-3 data [10] shows a sizable deviation from the other $e^{+}e^{-}$ data, which were used to derive the SM prediction for the muon $g-2$ in the dispersive approach taken in Ref. [7]. Thus, it is important to understand the hadronic contributions to the muon $g-2$ well within the SM and there is a need of improvements on the experimental data. Nonetheless, it becomes of interest to investigate what the impact of new physics on the stability issue is in the phenomenologically motivated models as discussed above and provide more information from the complementary tests of new physics.
In this article, we revisit a lepton portal model with an extra $U(1)^{\prime}$ gauge symmetry from this perspective, and perform an RG analysis to investigate on its stability. In this model, suggested by some of the authors [11], an $SU(2)_{L}$ singlet vector-like lepton charged under $U(1)^{\prime}$ is introduced. The vector-like lepton mixes with the muon through the vacuum expectation values (VEVs) of an extra Higgs doublet and a dark Higgs field, leading to a small seesaw mass for the muon. The muon $g-2$ can be explained thanks to extra one-loop corrections coming from the vector-lepton and the $Z^{\prime}$ boson, while the $W$ mass $m_{{}_{W}}$ can be increased, at the same time, through a tree-level mixing between the $Z$ boson and the $Z^{\prime}$ boson when the $Z^{\prime}$ is heavier than $Z$. As the muon $g-2$ anomaly favors the $U(1)^{\prime}$ breaking scale to be below about $200\,{\rm GeV}$, the extra gauge coupling required to increase $m_{{}_{W}}$ tends to be large [11], so that it can have a strong impact on the RG flow of the theory. There are previous analyses of RG equations and Higgs vacuum stability in the extended models with vector-like leptons and extended Higgs sectors with a motivation to explain the muon $g-2$ anomaly [12, 13]. In our work, making a complete study of the one-loop RG equations in our model, we aim to go beyond the limited discussion on loop corrections done in the previous work in Ref. [11]. As a result, we present new results on the stability and the Landau pole of the model in the parameter space favored by both the muon $g-2$ and $W$ boson mass anomalies. However, in view of the recent lattice results for the muon $g-2$, it is notable that our RG analysis is not limited to the parameter space for the muon $g-2$ anomaly.
@@ 3.1 SM anomalies: muon $g-2$ and $W$ boson mass
We comment on how flexible the parameter space is to explain the muon $g-2$ and $W$ boson mass anomalies. In the left plot of Fig. 1, we depict the one-loop correction to the muon $g-2$ as a function of $m_{{}_{Z^{\prime}}}$ in our model, with $v_{s}=v_{\phi}/\sqrt{2}=200\,{\rm GeV}$ and $150\,{\rm GeV}$. We set the parameters of the vector-like lepton sector to $M_{E}=1\,{\rm TeV}$, $\sin\theta_{L}=\sin\theta_{R}=0.01$, and choose $\sin\beta=0.18$ and $\lambda_{\phi}=0.1$ [11]. In the right plot of Fig. 1, we also show the tree-level correction to the $W$ boson mass due to the $Z$-$Z^{\prime}$ mixing as a function of $m_{{}_{Z^{\prime}}}$, with $v_{s}=v_{\phi}/\sqrt{2}=200\,{\rm GeV}$ and $150\,{\rm GeV}$, making the same choice for the other parameters as in the previous case. We note that in the case of $v_{s}=200\,{\rm GeV}$, new contributions to the muon $g-2$ in our model are in tension with the experimental value at $2\sigma$ for the SM prediction based on Ref. [7], but they can be consistent with the recent lattice results [9] at $2\sigma$. Thus, the parameter choices for Fig. 1 are compatible with the SM predictions appearing in the literature, so they will be taken for the later analysis on RG equations.
@@ 3.2 $Z^{\prime}$ signatures at the LHC
Let us now discuss the $Z^{\prime}$ production and its decay to the dilepton state. The $Z^{\prime}$ production cross sections (left) and $Z^{\prime}$ production cross sections multiplied by the $ee$ and $\mu\mu$ dilepton branching ratio (right) for $pp$ collision at $13$ TeV are depicted in Fig. 3 as a function of $m_{{}_{Z}^{\prime}}$. The opening of the $WW$ and $hZ$ channels is such that the dilepton production is suppressed strongly for large enough values of $m_{{}_{Z^{\prime}}}$, for e.g. it is less than $1$ fb for $m_{{}_{Z^{\prime}}}\gtrsim 300~{}$GeV.
With the results of ATLAS for dilepton final states at the $13$ TeV LHC run with integrated luminosity $L_{\mathrm{int}}=139$/fb [21], parametric regions with $m_{{}_{Z^{\prime}}}\gtrsim 300~{}$ GeV and $\sin\beta=0.18$ are allowed in our scenario. As we have seen, within this parametric region the $WW$ and $hZ$ final states are dominant, and these resonant boson productions are safe with respect to the $L_{\mathrm{int}}=139$/fb, $13$ TeV LHC results  [24, 22, 23, 25]. As the LHC constraints on dileptons are quite severe, dilepton final state produced from $Z^{\prime}$ boson resonance around $m_{{}_{Z^{\prime}}}\sim 300~{}$GeV should be targeted by the next LHC results. The same considerations hold for $WW$ final states for $m_{{}_{Z^{\prime}}}\gtrsim 300~{}$ GeV, with the latter being dominant within our scenario.
@@ 4.5 $Z^{\prime}$ mass dependence of the Landau pole
and calculating its Landau pole $\mu_{{}_{L}}=\mu_{{}_{\rm IR}}e^{{6\pi^{2}}/{7g^{2}_{{}_{Z^{\prime}}}(\mu_{{}_{\rm IR}})}}$ for the different IR values of $g_{{}_{Z^{\prime}}}$ in the two sets of the parameters in Table 4. The Landau pole of the full model is compared to $\mu_{{}_{L}}$ in Fig. 5, where it can be seen that the higher the value of $m_{{}_{Z^{\prime}}}$ ($g_{{}_{Z^{\prime}}}$), the closer the two are. It is easy to conclude from this plot that, for “high enough” (IR) values of $g_{{}_{Z^{\prime}}}$, the Landau pole is almost fully determined by the gauge coupling, that, through its contribution in their beta functions, causes the other couplings to diverge at slightly lower values. Lowering the (IR) value of $g_{{}_{Z^{\prime}}}$, the Landau pole of (45) becomes considerably higher than that of the full model. In this region of parameter space, the appearance of the Landau pole is a more “collective” phenomenon. All the couplings conspire together to generate it, and the determination of the pole scale from the RG equation of $g_{{}_{Z^{\prime}}}$ taken in isolation does not provide a good approximation to the pole scale of the full model any more.
Figure 6: Running $g_{{}_{Z^{\prime}}}$ coupling at one-loop and two-loop levels, in black solid and blue dashed lines, respectively. We considered two benchmark models with $(g_{{}_{Z^{\prime}}},\lambda_{E},y_{E})=(0.87,0.055,0.32)$ and $(1.15,0.073,0.32)$, on the left and right plots, respectively.
It is worthwhile to make comments on the effects of two-loop corrections in view of the low scale Landau pole, although a complete two-loop analysis in our model is beyond the scope of our current work. The RG equation for the extra gauge coupling at the two-loop level [34] is given by
$\mu\frac{d}{d\mu}g_{{}_{Z^{\prime}}}(\mu)=\frac{7}{12\pi^{2}}g_{{}_{Z^{\prime}}}^{3}+\frac{1}{(4\pi)^{4}}\bigg{(}256g_{{}_{Z^{\prime}}}^{5}-16g_{{}_{Z^{\prime}}}^{3}(\lambda_{E}^{2}+2y^{2}_{E})\bigg{)}.$ (46)
Then, ignoring the running of the extra Yukawa couplings, $\lambda_{E}$ and $y_{E}$, we get the approximate solution for the running extra gauge coupling at two-loops as
$g_{{}_{Z^{\prime}}}(\mu)\simeq\left[\frac{1}{g^{2}_{{}_{Z^{\prime}}}(\mu_{{}_{\rm IR}})}-\frac{98\log\left(\frac{\mu}{\mu_{{}_{\rm IR}}}\right)}{3(28\pi^{2}+3\lambda^{2}_{E}+6y^{2}_{E})}-\frac{96\log\left(\frac{g_{{}_{Z^{\prime}}}(\mu)}{g_{{}_{Z^{\prime}}}(\mu_{{}_{\rm IR}})}\right)}{28\pi^{2}+3\lambda^{2}_{E}+6y^{2}_{E}}\right]^{-1/2}.$ (47)
Then, we get the Landau pole modified to
$\mu_{{}_{L}}\simeq\mu_{{}_{\rm IR}}\,{\rm exp}\bigg{[}\frac{6\pi^{2}}{7g^{2}_{{}_{Z^{\prime}}}(\mu_{{}_{\rm IR}})}\Big{(}1+\frac{3\lambda^{2}_{E}}{28\pi^{2}}+\frac{3y^{2}_{E}}{14\pi^{2}}\Big{)}-\frac{144}{49}\log\left(\frac{g_{{}_{Z^{\prime}}}(\mu_{{}_{L}})}{g_{{}_{Z^{\prime}}}(\mu_{{}_{\rm IR}})}\right)\bigg{]}.$ (48)
Here, we can take the extra gauge coupling at the Landau pole in the correction term as $g_{{}_{Z^{\prime}}}(\mu_{{}_{L}})=4\pi$ for which perturbativity breaks down. For instance, for $\lambda_{E}=0.055-0.073$ and $y_{E}=0.32$ in our benchmark models in Section 4.3, we find that the scale of the Landau pole gets reduced sizably at two-loop order due to the two-loop gauge interactions, as compared to the one-loop results. In the left and right plots in Fig. 6, we also depict the running $g_{{}_{Z^{\prime}}}$ for the same benchmark models by the numerical analysis at one-loop and two-loop levels, in black solid and blue dashed lines, respectively. As a result, we find that the two-loop corrections are minor far away from the Landau pole, but they become important near the Landau pole. However, the very existence of the Landau pole at a low scale can be identified by the one-loop results, indicating the violation of perturbativity, so we can rely on the one-loop corrections in our RG analysis to extract the meaningful information on the Landau pole. Similar conclusions can be drawn for the perturbativity and stability of the other running couplings such as the Higgs quartic couplings in our model, but we postpone the detailed analysis with two-loop corrections to another work.